OpenStudy (anonymous):

The length of a rectangle is 3 times the width, and the perimeter is 22. Find the dimensions of the rectangle. {Width is , Length is

OpenStudy (anonymous):

so L = 3*W and from geometry we know that P = 2L + 2W = 22 plug and chug 22= 2(3W) + 2W = 8W ==> W= 22/8

OpenStudy (anonymous):

??

OpenStudy (aravindg):

perimeter=2(l+b) where l=length and b=breadth

OpenStudy (anonymous):



OpenStudy (aravindg):

here l=3b use the equation to get l and then you can get b

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

it kinda helps

OpenStudy (anonymous):

all i need is the l and w